# A Case Study on Misinformation Detection

### v3

@@ 4 Systematization of Literature
2. Antoniadis et al. [178] U A Acct metadata (U.i); syntax (A.i) Hurricane Sandy tweet dataset J48, RF, KNN, Bayes 0.79 (Avg. Prec.)
3. Assenmacher et al. [391] N A Propagation (N.i); topic det. (A.ii) Twitter (API) Clustering not reported
@@
4. Buntain et al. [50] N U A Time (N.ii); acct metadata (U.i); sentiment (A.i) CREDBANK, Buzzfeed RF 0.65 (Acc.)
@@
5. Castillo et al. [253] U A Syntax (A.i); user behavior (U.ii) Twitter Monitor events SVM, DT 0.874 (P)
6. Chen et al. [183] N U A Social graph (N.iii); syntax (A.i); user behavior (U.ii) Weibo RNN 0.92 (Acc.)
7. Guo et al. [272] N C Prop. (N.i); semantics (C.i) Twitter, Weibo LSTM 0.9 (Acc.)
8. Jin et al. [263] N A Propagation (N.i); stance (A.iii) Sina Weibo posts Clustering 0.84 (Acc.)
9. Liu et al. [40] N Propagation (N.i) Weibo, Twitter15, Twitter16 RNN, CNN 0.897 (Acc.)
10. Ma et al. [30] N Propagation (N.i) Kochina, Ma, Shu Twitter datasets RNN/BiLSTM 0.75 (Acc.)
11. Magelinski et al. [262] N U Prop. (N.i); timing (N.ii); user behavior (U.ii) Twitter (API) – n/a
12. Nguyen et al. [69] N A Prop. (N.i); semantics (A.i) Twitter, Weibo, PHEME [146] SVM, RNN, DT 0.970 (Acc.)
13. Pacheco et al. [92] U N Account metadata (U.i); timing (N.i) Twitter (API) Clustering 0.8+ (Prec.)
14. Ratkiewicz et al. [388] N A U Prop. (N.i); keywords (A.ii); user behavior (U.ii) Twitter (API) AdaBoost, SVM 0.96 (Acc.)
15. Sharma et al. [158] N A U Prop. (N.i); keywords (A.ii); user behavior (U.ii) Twitter (API); Twitter IRA trolls list GMM, Kmeans, NN 0.94 (AUC)
16. Tschiatschek et al. [230] N U Prop. (N.i); user rep. (U.i) Facebook dataset Bayes inf. n/a
17. Zeng et al. [133] N A Prop. (N.i); stance (A.iii) 4.3K tweets (from API) Logit, NB, RF 0.88 (Acc.)
@@
1. Asr et al. [287] W A Source rep. (W.i); Syntax (A.i) BuzzfeedUSE, Snopes, Rashkin, Rubin CNN, SVM, NB not reported
@@
2. Baly et al. [43] W A N Source rep. (W.i); Site infra. (W.ii); (A.i) MediaBiasFactCheck[311] SVM 0.66 (Acc.)
@@
3. Baly et al. [316] W A N Source rep. (W.i); Site infra. (W.ii); (A.i) MediaBiasFactCheck[311] SVM 0.7152 (Acc.)
4. Castelo et al. [302] W A Site infra. (W.ii); syntax (A.i) Celebrity, US-Election2016 SVM, kNN, RF 0.86 (Acc.)
5. Chen et al. [109] W A Hosting infra. (URL) (W.ii); syntax (A.i) PoliticalFakeNews Clustering 0.97 (AUC)
6. Hounsel et al. [1] W Site infra. (W.i) FactCheck, Snopes, PolitiFact, Buzzfeed RF 0.98 (AUC)
7. Ribeiro et al. [254] W U Site infra. (W.i); User demog. (U.i) Facebook API Graph search 0.97 (PCC)
@@
TABLE I: Focus corpus by scope and target. Coding of a focus set of 87 papers, sorted by information scope. Values in parentheses in “Target” field correspond to highlighted subcategories presented in Section 4 (e.g., “C.i” denotes target (i) in “Claims,” Section 4.1). If authors present evaluation results for multiple models, we underline the most performant model and record its corresponding performance score.
@@ 4.1 Claims
About 70% of paper authors within this scope cite the 2016 U.S. presidential election as motivation for the development of their methods [84, 130, 297]; about 30% cite COVID-19 misinformation [128, 121, 142]. All works mention the speed and volume of social media misinformation in particular [108]. 30% cite the relatively slow pace of manual fact-checking [22, 130, 108]—and the need for faster, automated approaches—as motivation. Claim-scoped papers form 15% of our corpus.
@@
Detection targets. Across all information scopes, we find that claim-scoped detection methods are most consistent in their attempts to verify semantic contents of text statements. This is done in the following ways: (C.i) distance calculations on semantic embeddings to perform textual entailment or stance detection [77, 48]; and (C.ii) search on a knowledge graph topology [22, 56] to determine if these reference sources corroborate or refute the claim to be checked. A small class of approaches explicitly detect (C.iii)checkworthiness,”  [^9] We include these works in our corpus because they do take topic and source credibility into consideration in the process of ranking checkability. and are intended to surface checkable statements to human fact-checkers for manual verification [108, 56, 130]. Targets such as author credibility and language cues are proxy targets for the presence of misinformation: signals which may not be sufficient for determining text veracity in isolation, but which are indicative of (lack of) veracity by association with an external heuristic. About 60% of papers within the full corpus at this scope employ non-semantic targets, including the (A.i) syntactic and/or stylometric qualities of text [238], (U.i) source reputation [383], or (U.ii) contextual indicators, such as commenter responses [106], to classify social media posts.
@@
Datasets. Claim-scoped papers in our corpus that propose to perform fact verification (2.i) rely on outdated existing datasets of labeled statements in order to establish ground truth. LIAR [35] [^10] We note that the LIAR dataset is actually a collection of 27.8K labeled PolitiFact statements—so, in a sense, PolitiFact is the dominant data source. and PolitiFact [299] were the most popular datasets among claim-scoped works, and, taken together, were used by approximately half of all papers within this scope [76, 123]. LIAR is a static political news dataset that was published in 2017; on average, papers that cite LIAR were published two years after LIAR’s release. Topic detection and word frequency models trained on LIAR are likely ineffective in contemporary fact-checking contexts, and for non-political subject matter [84, 383]. [^11] Choice of ground truth site or labeled dataset can significantly influence the outcome of analysis: Bozarth et al. found that perceived prevalence of misinformation in a corpus of 2016 election news varied from 2% to 40%, depending on choice of ground-truth reference website [20]. Additionally, misinformation taxonomies across reference sites are inconsistent: PolitiFact employs a six-point labeling scale (pants-on-fire; false; barely-true; half-true; mostly-true; true), FEVER employs a three-point scale (supported; refuted; notenoughinfo), and GossipCop employs an eleven-point scale (ratings from 0 to 10). This divergence is likely a symptom of definitional issues (Section 3).
@@
Model selection. Model choice follows target choice for claim-scoped methods: for methods that pre-construct knowledge graphs or other reference databases, models perform some form of (3.i) shortest-path search on the KG topology [56, 22], and approximate logical inference via transitive closure on graph edges. For methods that perform information retrieval at query time—e.g., to match corroborating sources to a claim to be checked—model training generally follows conversion of the text statement to a bag-of-words or TF-IDF embedding; choice of model is highly variable [236, 207], and does not appear to be predictive of performance. For methods that perform detection of misinformative posts on social media, (3.v) stacked ensemble classifiers are a common approach to incorporating multiple feature modalities.
@@
Feature selection. At the level of single claims, semantic feature analysis is limited to the (4.i) identification of structured statements as a precursor to knowledge graph (KG) construction. These claims take the form of subject-predicate-object (SPO) statements (e.g., “I like pie”) [22, 18, 56]. Detection versatility is determined by the size of the source dataset and the granularity of the relationships encoded by graph edges [56]. Supervised methods that detect linguistic cues employ (4.i) hand-crafted word or topic lists [108]). Authors employing supervised methods claim that their approach permits highly customized targeting of specific rumoring narratives [108], though this also assumes that the method developer has prior knowledge of the contents of test data; authors employing unsupervised methods claim that their approaches detect contextual and language features that cannot be easily extracted by common features such as word frequency or sentiment [76]. Methods detecting social media data consider (4.ii) network and (4.iii) user interaction features (post likes, shares, comments) [185, 294, 277].
@@
Evaluation. Though a majority of claim-scoped methods cite the speed of social media misinformation as motivation, (5.i) only one method within this scope reported results from testing in real time [54]. KG-based methods are (5.ii) non-generalizable by design: the approaches we survey require structured inputs for graph construction, and rely on published datasets for ground truth [22, 56]. Though this approach permits semantic verification of statements, it is difficult to perform iterative updates to source databases in real-time, particularly in the types of online settings where claim-checking might be most usefully deployed (e.g., during breaking news events, where no source of ground truth is immediately available). Additionally, we observe that a majority of works at this scope do not test on novel or out-of-domain data; we discuss overfitting issues in greater detail in the next section.
Takeaways: 1 The efficacy of claim-scoped methods is completely determined by the depth and breadth of coverage conferred by a reference database. 2 Datasets are frequently out-of-date, and taxonomies are inconsistent. 3 Though some inference is possible via transitive closure on knowledge graph edges, this capability is, in general, limited. 4 Knowledge-graph-based methods require structured inputs, which might not be readily available in a breaking news setting, or in scenarios where ground truth references are not yet available. 5 Few authors test in real-time, despite citing the slow pace of manual fact-checking as motivation for their work.
4.2 Articles
We consider all news-oriented writing of length 100 words or greater to fall within this scope. 25% of papers within this scope cite growing distrust of mainstream news media outlets as motivation for their methods, which promise to deliver fast labeling of news stories that appear on social media [174]. Text-based credibility classifiers have been shown to have limited efficacy, however: while unsupervised approaches can identify bias with high accuracy, this performance degrades significantly in misinformation and credibility classification tasks [395, 17]. This scope comprises 24% of our full paper corpus.
@@ 4.2 Articles
Detection targets. In contrast to single claims, full-length news articles have sufficient text contents to make semantic verification difficult—and certain off-the-shelf NLP approaches practicable. These approaches are distinct from direct claim verification and qualify as proxy detection methods: For instance, Bhutani et al. associate strongly negative sentiment with the presence of false information [236], and Horne et al. find that satire and misinformation share stylistic similarities [24]. All article-scoped methods target proxy signals, and adopt at least one of the following three approaches to detection: (A.i) NLP analysis of article contents to identify language features particular to writing styles heuristically associated with misinformation (genre detection, sentiment analysis) [324]; and (A.ii) analysis of article contents and headlines to identify potentially clickbait-y titles or discussion related to known misinformation narratives (topic detection) [163, 108]. Respectively, these approaches 1) simultaneously assume and detect a heuristic (e.g., strong emotion indicating the presence of misinformation); and 2) assume prior knowledge of rumoring topics.
@@
Datasets. While well-annotated, current datasets are in short supply across all information scopes, this deficit is particularly glaring at the article scope. This is due in large part to definitional ambiguities that prevent fine-grained labeling of longer texts for classifier training. As a result, 44% of paper authors within this scope (2.i) use public datasets released years prior to the start of their research [68, 35, 24, 41]. These datasets (LIAR [35], Buzzfeed-Webis [17], and PolitiFact [299]) include news links, speaker credibility scores, and other metadata that (2.ii) constitute serious sources of leakage for methods that use contextual features to infer true/false labels. The remaining 56% of authors curate their own article corpora by asking crowdworkers to generate misinformative text [63], selectively editing true news articles (e.g., via verb inversion or noun replacement) [31], or compiling articles from authoritative news sources and known satire sites [240]. These data curation techniques (2.iii) introduce additional dependencies and shortcuts to textual datasets for which such variables are already difficult to control. Style [17] and genre [89], for instance, are emergent qualities of writing that cannot be easily marginalized out of a text embedding.
@@
Model selection. Among papers that report testing with multiple models—including (3.ii) classical ML models [214], (3.iii) unsupervised NN models [68], and (3.v) stacked ensemble classifiers [37]—there is no clear correlation between model choice and actual performance. We note that, in instances where authors test on two- and multi (i.e., $>$ 2)-way classification tasks, performance declines sharply in the latter case [212]. In those instances, reported performance scores are for two-way tasks. For this reason, as well, classical ML models (logit, SVM) oftentimes appear to be most performant.
@@
Feature selection. Among supervised methods that disclose their feature sets, we find that (4.i) word frequency, sentiment, and genre were among the most commonly used features, and were collectively employed by 80% of works within this scope; these features can also be sources of dependency-induced noise. It is difficult to quantify the impact of dependencies related to voice, house style, and source on classifier performance, particularly in the case of unsupervised learning methods, which comprise 59% of methods at this scope. We evaluate an unsupervised learning method (and consider its performance in light of possible style-related dependencies) in our replication analysis of Nasir et al. (Section 5.1) [141].
@@
Evaluation. Misinformation detection methods scoped to full texts risk overfitting to single topics: 42% of authors select (5.ii) one or more narratives of interest (e.g., the 2016 presidential election, the Boston Marathon bombing), train a classifier on these topics, then test this classifier on a different set of texts that discuss the same topic [276, 216, 141]. This approach, while valid for evaluating classifier performance on closed datasets, lacks ecological validity for the use cases that authors claim that their methods will address: Rapid topic identification and high-quality annotation of relevant articles are generally unavailable in breaking news scenarios on social media platforms [396]. 60% of authors at this scope compare the performance of their detection method to other published approaches or ML models, but (5.ii) neglect to test on novel datasets. These methods do well when tested on in-domain texts, and in comparison to a selection of older ML models; many report accuracy well above 80% [73, 111]. Only one paper within the article scope tested in an adversarial setting: Its authors found that, while stylometry-based misinformation detection had an accuracy rate greater than 80% on routine tasks, this score dropped to about 50% in adversarial cases [212]. We demonstrate such a performance dropoff in our replication analysis (Section 5.1).
Takeaways: 1 In the absence of semantic definitions of misinformation, proxy detection targets are common but easy to evade. 2 Well-labeled datasets are rare; those datasets that are available are at least several years old at time of writing. 3 Unsupervised methods show marginal improvements over classical ML models in some cases; it is unclear if these improvements are 1) significant or 2) sustainable across different datasets. 4 Detection methods at the article scope are uniquely susceptible to text-based dependencies that are difficult to control for. 5 Inflated performance scores can often be attributed to testing on same-topic news articles.
4.3 Users
Evidence of foreign interference during the 2016 U.S. presidential election triggered a resurgent interest in malicious account detection [366, 361]. As such, 90% of security- and social-science-oriented works that we include within this scope (a dozen papers) explicitly discuss Russian trolls or other influence operations conducted by nation state actors and train classifiers on published lists of such accounts [361, 362, 360, 364, 365]. Account-scoped papers formed approximately 15% of our corpus (40 papers).

### v4

@@ 4 Systematization of Literature
2. Antoniadis et al. [antoniadis_model_2015] U A Acct metadata (U.i); syntax (A.i) Hurricane Sandy tweet dataset J48, RF, KNN, Bayes 0.79 (Avg. Prec.)
3. Assenmacher et al. [assenmacher2020two] N A Propagation (N.i); topic det. (A.ii) Twitter (API) Clustering not reported
@@
4. Buntain et al. [buntain_automatically_2017] N U A Time (N.ii); acct metadata (U.i); sentiment (A.i) CREDBANK, Buzzfeed RF 0.65 (Acc.)
@@
5. Castillo et al. [castillo_information_2011] U A Syntax (A.i); user behavior (U.ii) Twitter Monitor events SVM, DT 0.874 (P)
@@
1. Asr et al. [torabi_asr_data_2018] W A Source rep. (W.i); Syntax (A.i) BuzzfeedUSE, Snopes, Rashkin, Rubin CNN, SVM, NB not reported
@@
2. Baly et al. [baly_predicting_nodate] W A N Source rep. (W.i); Site infra. (W.ii); (A.i) MediaBiasFactCheck[mediabiasfactheck] SVM 0.66 (Acc.)
@@
3. Baly et al. [baly-etal-2020-written] W A N Source rep. (W.i); Site infra. (W.ii); (A.i) MediaBiasFactCheck[mediabiasfactheck] SVM 0.7152 (Acc.)
4. Castelo et al. [castelo_agnostic_2019] W A Site infra. (W.ii); syntax (A.i) Celebrity, US-Election2016 SVM, kNN, RF 0.86 (Acc.)
5. Chen et al. [chen_proactive_2020] W A Hosting infra. (URL) (W.ii); syntax (A.i) PoliticalFakeNews Clustering 0.97 (AUC)
@@
Table 2: Focus corpus by scope and target. Coding of 25 papers in our focus set, sorted by information scope. Codes for the full focus set appear in Table 4. Values in parentheses in “Target” field correspond to highlighted subcategories presented in Section 4 (e.g., “C.i” denotes target (i) in “Claims,”  Section 4.1). If authors present evaluation results for multiple models, we underline the most performant model and record its corresponding performance score.
@@ 4.1 Claims
About 70% of paper authors within this scope cite the 2016 U.S. presidential election as motivation for the development of their methods [debnath_hierarchical_2020, hassan_data_2014, katsaros_which_2019]; about 30% cite COVID-19 misinformation [kou_hc-covid_2022, roitero_can_2021, glockner_missing_2022]. All works mention the speed and volume of social media misinformation in particular [kartal_too_2020]. 30% cite the relatively slow pace of manual fact-checking [ciampaglia-wiki-2015, hassan_data_2014, kartal_too_2020]—and the need for faster, automated approaches—as motivation. Claim-scoped papers form 15% of our corpus.
@@
Detection targets. Across all information scopes, we find that claim-scoped detection methods are most consistent in their attempts to verify semantic contents of text statements. This is done in the following ways: (C.i) distance calculations on semantic embeddings to perform textual entailment or stance detection [saikh_novel_2019, borges_combining_2018]; and (C.ii) search on a knowledge graph topology [ciampaglia-wiki-2015, shiralkar_finding_2017] to determine if these reference sources corroborate or refute the claim to be checked. A small class of approaches explicitly detect (C.iii)checkworthiness,”  [^9] We include these works in our corpus because they do take topic and source credibility into consideration in the process of ranking checkability. and are intended to surface checkable statements to human fact-checkers for manual verification [kartal_too_2020, shiralkar_finding_2017, hassan_data_2014].
Targets such as author credibility and language cues are proxy targets for the presence of misinformation: signals which may not be sufficient for determining text veracity in isolation, but which are indicative of (lack of) veracity by association with an external heuristic. About 60% of papers within the full corpus at this scope employ non-semantic targets, including the (A.i) syntactic and/or stylometric qualities of text [jadhav_fake_2019], (U.i) source reputation [rasool_multi-label_2019], or (U.ii) contextual indicators, such as commenter responses [tian_early_2020], to classify social media posts.
4
Feature selection. At the level of single claims, semantic feature analysis is limited to the (4.i) identification of structured statements as a precursor to knowledge graph (KG) construction. These claims take the form of subject-predicate-object (SPO) statements (e.g., “I like pie”) [ciampaglia-wiki-2015, magdy-stat-2010, shiralkar_finding_2017]. Detection versatility is determined by the size of the source dataset and the granularity of the relationships encoded by graph edges [shiralkar_finding_2017]. Supervised methods that detect linguistic cues employ (4.i) hand-crafted word or topic lists [kartal_too_2020]). Authors employing supervised methods claim that their approach permits highly customized targeting of specific rumoring narratives [kartal_too_2020], though this also assumes that the method developer has prior knowledge of the contents of test data; authors employing unsupervised methods claim that their approaches detect contextual and language features that cannot be easily extracted by common features such as word frequency or sentiment [wang_learning_2019]. Methods detecting social media data consider (4.ii) network and (4.iii) user interaction features (post likes, shares, comments) [paudel_lambretta_2023, gupta_tweetcred_2015, cui_same_2019].
5
Evaluation. Though a majority of claim-scoped methods cite the speed of social media misinformation as motivation, (5.i) only one method within this scope reported results from testing in real time [hassan_toward_2017]. KG-based methods are (5.ii) non-generalizable by design: the approaches we survey require structured inputs for graph construction, and rely on published datasets for ground truth [ciampaglia-wiki-2015, shiralkar_finding_2017]. Though this approach permits semantic verification of statements, it is difficult to perform iterative updates to source databases in real-time, particularly in the types of online settings where claim-checking might be most usefully deployed (e.g., during breaking news events, where no source of ground truth is immediately available). We observe that a majority of works at this scope do not test on novel or out-of-domain data; we discuss overfitting issues in greater detail in the next section.
**Takeaways: 1 The efficacy of claim-scoped methods is completely determined by the depth and breadth of coverage conferred by a reference database. 2 Datasets are frequently out-of-date, and taxonomies are inconsistent. 3 Though some inference is possible via transitive closure on knowledge graph edges, this capability is, in general, limited. 4 Knowledge-graph-based methods require structured inputs, which might not be readily available in a breaking news setting, or in scenarios where ground truth references are not yet available. 5 Few authors test in real-time, despite citing the slow pace of manual fact-checking as motivation.
4.2 Articles
We consider all news-oriented writing of length 100 words or greater to fall within this scope. 25% of papers within this scope cite growing distrust of mainstream news media outlets as motivation for their methods, which promise to deliver fast labeling of news stories that appear on social media [rubin_deception_2015]. Text-based credibility classifiers have been shown to have limited efficacy, however: while unsupervised approaches can identify bias with high accuracy, this performance degrades in misinformation and credibility classification tasks [fairbanks2018credibility, potthast-etal-2018-stylometric]. This scope comprises 24% of our paper corpus.
1
Detection targets. In contrast to single claims, full-length news articles have sufficient text contents to make semantic verification difficult—and certain off-the-shelf NLP approaches practicable. These approaches are distinct from direct claim verification and qualify as proxy detection methods: For instance, Bhutani et al. associate strongly negative sentiment with the presence of false information [bhutani_fake_2019], and Horne et al. find that satire and misinformation share stylistic similarities [horne_this_2017]. All article-scoped methods target proxy signals, and adopt at least one of the following three approaches to detection: (A.i) NLP analysis of article contents to identify language features particular to writing styles heuristically associated with misinformation (genre detection, sentiment analysis) [mihalcea-strapparava-2009-lies]; and (A.ii) analysis of article contents and headlines to identify potentially clickbait-y titles or discussion related to known misinformation narratives (topic detection) [zhang_fake_2018, kartal_too_2020]. Respectively, these approaches 1) simultaneously assume and detect a heuristic (e.g., strong emotion indicating the presence of misinformation); and 2) assume prior knowledge of rumoring topics.
@@ 4.2 Articles
Datasets. While well-annotated, current datasets are in short supply across all information scopes, this deficit is particularly glaring at the article scope. This is due in large part to definitional ambiguities that prevent fine-grained labeling of longer texts for classifier training. As a result, 44% of paper authors within this scope (2.i) use public datasets released years prior to the start of their research [brasoveanu_semantic_2019, wang_liar_2017, horne_this_2017, shu_beyond_2019]. These datasets (LIAR [wang_liar_2017], Buzzfeed-Webis [potthast-etal-2018-stylometric], and PolitiFact [politifact]) include news links, speaker credibility scores, and other metadata that (2.ii) constitute serious sources of leakage for methods that use contextual features to infer true/false labels. The remaining 56% of authors curate their own article corpora by asking crowdworkers to generate misinformative text [perez-rosas_automatic_2017], selectively editing true news articles (e.g., via verb inversion or noun replacement) [mihalcea_lie_2009], or compiling articles from authoritative news sources and known satire sites [rubin_fake_2016]. These data curation techniques (2.iii) introduce additional dependencies and shortcuts to textual datasets for which such variables are already difficult to control. Style [potthast-etal-2018-stylometric] and genre [hosseinimotlagh_unsupervised_nodate], for instance, are emergent qualities of writing that cannot be easily marginalized out of a text embedding.
@@
Model selection. Among papers that report testing with multiple models—including (3.ii) classical ML models [ahmed_detecting_2018], (3.iii) unsupervised NN models [brasoveanu_semantic_2019], and (3.v) stacked ensemble classifiers [ruchansky_csi_2017]—there is no clear correlation between model choice and actual performance. We note that, in instances where authors test on two- and multi (i.e., $>$ 2)-way classification tasks, performance declines sharply in the latter case [afroz_detecting_2012]. In those instances, reported performance scores are for two-way tasks. For this reason, as well, classical ML models (logit, SVM) oftentimes appear to be most performant.
@@
Feature selection. Among supervised methods that disclose their feature sets, we find that (4.i) word frequency, sentiment, and genre were among the most commonly used features, and were collectively employed by 80% of works within this scope; these features can also be sources of dependency-induced noise. It is difficult to quantify the impact of dependencies related to voice, house style, and source on classifier performance, particularly in the case of unsupervised learning methods, which comprise 59% of methods at this scope. We evaluate an unsupervised learning method (and consider its performance in light of possible style-related dependencies) in our replication analysis of Nasir et al. (Section 5.1) [nasir_fake_2021].
@@
Evaluation. Misinformation detection methods scoped to full texts risk overfitting to single topics: 42% of authors select (5.ii) one or more narratives of interest (e.g., the 2016 presidential election, the Boston Marathon bombing), train a classifier on these topics, then test this classifier on a different set of texts that discuss the same topic [noauthor_rumors_2014, jin_detection_2017, nasir_fake_2021]. This approach, while valid for evaluating classifier performance on closed datasets, lacks ecological validity for the use cases that authors claim that their methods will address: Rapid topic identification and high-quality annotation of relevant articles are generally unavailable in breaking news scenarios on social media platforms [shu2020detecting]. 60% of authors at this scope compare the performance of their detection method to other published approaches or ML models, but (5.ii) neglect to test on novel datasets. These methods do well when tested on in-domain texts, and in comparison to a selection of older ML models; many report accuracy well above 80% [zhang_detecting_2019, silva_embracing_2021]. Only one paper within the article scope tested in an adversarial setting: Its authors found that, while stylometry-based misinformation detection had an accuracy rate greater than 80% on routine tasks, this score dropped to about 50% in adversarial cases [afroz_detecting_2012]. We demonstrate such a dropoff in Section 5.1.
**Takeaways: 1 In the absence of semantic definitions of misinformation, proxy detection targets are common but easy to evade. 2 Well-labeled datasets are rare; those datasets that are available are at least several years old at time of writing. 3 Unsupervised methods show marginal improvements over classical ML models in some cases; it is unclear if these improvements are 1) significant or 2) sustainable across different datasets. 4 Detection methods at the article scope are uniquely susceptible to text-based dependencies that are difficult to control for. 5 Inflated performance scores can often be attributed to testing on same-topic news articles.
4.3 Users
Evidence of foreign interference during the 2016 U.S. presidential election triggered a resurgent interest in malicious account detection [ezzeddine2023exposing, saeed2022trollmagnifier]. As such, 90% of security- and social-science-oriented works that we include within this scope (a dozen papers) explicitly discuss Russian trolls or other influence operations conducted by nation state actors and train classifiers on published lists of such accounts [saeed2022trollmagnifier, zannettou2019disinformation, mirza2023tactics, lukito2020coordinating, keller2020political]. Account-scoped papers formed approximately 15% of our corpus (40 papers).
@@ 4.3 Users
Detection targets. Papers within this scope target source reputation, and (U.i) inspect account metadata, such bios, account age, and profile images; or distinguish suspicious accounts by (U.ii) a single user’s social behaviors, such as their comments on posts (n.b. this target is distinct from (N.i)). The security literature we review discusses trolls and bots deployed for astroturfing, misinformation campaigns, and IOs [saeed2022trollmagnifier, cresci_bots_2020, chu2012detecting, zhang2016rise, alizadeh2020content]. In the absence of rigorous definitions of these account types, however, actual detection targets are tautological: A troll or bot is an account that exhibits troll- or bot-like behavior, or that interacts with confirmed troll or bot accounts [saeed2022trollmagnifier, cresci_bots_2020].
@@
Datasets. All troll and bot account detection works we reviewed relied on published lists of “known” troll accounts for model training, but (2.iv) neglected to mention the heavily manual investigation required to produce these original lists [silverman_timberg_kao_merrill_2022, linvill2020troll]. Researchers who compiled some of these account lists, including a set of several hundred Twitter accounts with possible links to a known Russian troll farm (the Internet Research Agency, or IRA) manually examined suspicious accounts and tweet contents in order to produce detailed account and content taxonomies; notably, these classifications required external intelligence about account activity that was not published alongside account lists [linvill2020troll]. Two well-cited troll lists compiled by the U.S. government, comprising thousands of suspicious Twitter and Facebook accounts, were curated using proprietary non-public information [exhibit_b].
@@
Model selection. Detection proceeds via classifier training on a list of “known” suspicious accounts and application of this classifier to a dataset of novel accounts [saeed2022trollmagnifier, lukito2020coordinating]. We note that, regardless of model choice, if classifier training data and feature selection reflect a heuristic about suspicious behavior, the resulting classifier will simply learn this heuristic: The methods we review can be used to detect accounts whose behaviors conform to heuristic assumptions, but cannot be used to surface novel malicious behaviors, and are not resistant to attacks or evasion [wang2014man]; we explore this further in our replication analysis of Saeed et al. (Section 5.2) [saeed2022trollmagnifier]. A subset of methods at this scope and the network scope formulate detection as a (3.iv) graph cut or influence maximization problem, and describe approaches to identifying optimal cuts for isolating suspicious accounts [zhang_misinformation_2016, cao_aiding_2012].
@@
Feature selection. Methods that define suspicious accounts by intrinsic properties of these accounts (e.g., user handles and profile images) target the (4.iii) semantics of this account metadata, and detect evidence of manipulation in (e.g.) image metadata and bios; or text outputs, such as posts and links [helmstetter-supervised-2018, sansonetti_unreliable_2020]. Methods that define suspicious accounts by account activity target (4.ii) networked behaviors, such as liking and resharing statistics [pacheco_uncovering_2020]. Feature sets for some methods in the first category include demographic data for users, such as inferred political party affiliation or race [ribeiro_media_2018] (these nth order assumptions are dangerous to make [freelon_black_2020]; see our discussion about proxy signals, in Section 4.2).
@@
Evaluation. We observe accuracy scores above 80% for (5.ii) confirmatory detection of like accounts for all methods that reference a seed list of known trolls [saeed2022trollmagnifier, addawood_linguistic_2019, ezzeddine2023exposing]; de novo detection of behaviors not represented within training data is not possible, by the self-admission of 10% of authors within this scope [saeed2022trollmagnifier, stone2009botnet]. As we discuss further in the next subsection (Networks), the increasingly hybrid nature of IOs requires more nuanced taxonomies for classification: extent of coordination, rather than existence, might be a more appropriate measure of possible manipulation.  [^10] In a blog post on bot detection from 2021, Twitter “debunked” four common heuristics commonly used to identify bot accounts, including several detected by methods in our corpus, and described a “forensic team of investigators” who manually verify bot-ness of suspect accounts [Thread_2021]. Some authors of bot detection methods acknowledge that their approaches are (5.iii) trivially easy to evade if account holders 1) avoid interacting with known suspicious accounts or 2) vary their account identity and posting semantics [danezis2009sybilinfer, cresci_bots_2020].
**Takeaways: 1 Targets are frequently tautological. 2 Hand-annotated training datasets are the result of intensive fact-finding on the part of human researchers, and often require information that is not publicly available. 3 Classifiers can only detect accounts resembling those in seed lists. 4 Features that attempt to infer user credibility from demographic information risk reinforcing existing biases. 5 Current methods cannot detect novel malicious behaviors.
4.4 Networks
Within the security literature, a growing awareness of hybrid networks, which employ a combination of automated and manual approaches to disseminate content, has encouraged a turn toward network-based bot detection methods, and away from detection of individual accounts [cresci_bots_2020, grimme2018changing]. We observe a parallel turn toward network-based methods in AI, ML, and NLP venues as a result of growing recognition of overfitting and generalizability issues in purely text-based detection methods (see Articles) [castelo_agnostic_2019]. The common assumption, across disciplines, is that coordinated networks leave more detectable evidence of manipulation than do individual accounts, and that these footprints should be identifiable regardless of attack type or rumoring topic [castelo_agnostic_2019, milajerdi2019holmes]. Network-scoped methods, including relevant security literature, form 20% of our corpus.